An electric current of 104.0 mA transports 1.00 kC of charge. Calculate the time this took. Be sure your answer has the correct unit symbol and the correct number of significant digits. ×10
Answer #1

Q = 1.00 kC

= 1.00*10^3 C

I = 104.0 mA

= 0.1040 A

use:

I = Q/t

0.1040 A = (1.00*10^3 C) / t

t = 9.62*10^3 s

Answer: 9.62*10^3 s